OpenStudy (anonymous):

Solve for x please:

OpenStudy (anonymous):

\[3^{3x - 1} = 7^{x +2}\] small enough?

OpenStudy (dumbcow):

logs to the rescue

OpenStudy (dumbcow):

use property: \[\ln (a^{n}) = n*\ln(a)\]

OpenStudy (dumbcow):

did that help ... should i continue?

OpenStudy (anonymous):

Please continue. From: \[27^x = 147*7^x\]

OpenStudy (anonymous):

I got it actually. Answer is just so different from wolfram.

OpenStudy (dumbcow):

ok great :)

OpenStudy (dumbcow):

yeah wolfram may give in diff form ... scroll down to get alternate forms

OpenStudy (anonymous):

\[x = \log_{27/7} (147)\] but wolfram is: \[\frac{ \log(3) + 2\log(7) }{ 3\log(3) - \log(7) }\]

OpenStudy (dumbcow):

wolfram will use natural log by default

OpenStudy (dumbcow):

log with diff base is usually not standard form of answer....because you cant check it with calculator

OpenStudy (anonymous):

you can check it with a calculator though :) log_b(x) = log(x)/log(b)

OpenStudy (dumbcow):

:) so standard answer would be \[x = \frac{\ln 147}{\ln 27/7}\]